A random sample of 22 chocolate chip cookies were chosen from a bakery. The cookies were found to have an average of 19.094 chocolate chips with a standard deviation of 2.68. Find the 95% confidence interval for the true mean. The lower bound of the interval is _______
Answer #1

df = n - 1 = 22 - 1 = 21

From T table,

t critical value at 0.05 significance level with 21 df = 2.080

95% confidence interval for is

- t * S / sqrt(n) < < + t * S / sqrt(n)

19.094 - 2.080 * 2.68 / sqrt(22) < < 19.094 + 2.080 * 2.68 / sqrt(22)

17.906 < < 20.282

95% C I is ( 17.906 , 20.282 )

Lower bound = 17.906

Upper bound = 20.282